Fermi National Laboratory
 

Accelerator Update for February 18, 2009 - February 20, 2009

 

Wednesday Feb 18

The day shift began with Operations monitoring the Tevatron (TeV) store 6812 (75.03E30) and the Recycler stash (323.2E10), with the Antiproton Source stacking (28.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

At 8:14 AM, a Booster RF station (BRF6) tripped off for the first of three times in 11 minutes. An expert investigated after the last trip.

At 8:38 AM, Operations reported a problem with Booster RF sparking on the first pulse after a Neutron Therapy Facility (NTF) pulse. Experts investigated and found that the feed forward signal for LRF4 was learning in on the wrong pulse. This was corrected.

Operations terminated store 6812 at 10:25 AM.

Operations put the TeV into shot setup at 10:37 AM.

Operations reported at 11:47 AM that they had received permission to send beam to the Meson MTest experiment T971, the Pixel Telescope.

Operations established store 6813 at 12:11 PM with an initial luminosity of 325.53E30.

NTF reported at 12:33 PM that they had completed treating their last patient for the day.

Operations established beam to MTest at 1:45 PM.

A MI expert began a MI-8 multiwire study at 2:25 PM.

Another MI expert began study at 2:43 PM that lasted about three hours.

TeV experts began an alpha scan study at 3 PM that lasted about 3 hours.

Operations reported at 4:04 PM that the Recycler had stashed antiprotons six times during the day shift.

Operations disabled the MTest beam at 6:01 PM.

At 10:48 PM, Operations reported trouble with transfers of antiprotons from Pbar to the Recycler. Accumulator devices were being set improperly and causing emittance growth. An expert helped with the investigation. Eventually, Operations rebooted the Accumulator PLC and this fixed the problems.

Operations reported at midnight that the Recycler had stashed eight times during the evening shift.

Thursday Feb 19

The midnight shift began with Operations monitoring store 6813 (93.08E30) and the stash (303.4E10), with Pbar stacking (18.4E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 6813 at 2:53 AM. Operations put the TeV into shot setup.

At 3:27 AM, Operations reported that shortly after putting the TeV into shot setup, a TeV magnet (C:AQ7) quenched; its output current jumped by 30 amps. Operations called in an expert.

A TeV power supply expert swapped the AQ7 regulator at 4:56 AM.

Operations put the TeV through a dry squeeze at 5:05 AM.

Operations put the TeV through a wet squeeze at 5:46 AM.

Operations put the TeV into shot setup at 6:06 AM.

Operations prepared Pbar for access at 7:06 AM.

Operations prepared the Pelletron for access at 7:13 AM.

Operations established store 6816 at 7:23 AM with an initial luminosity of 323.28E30.

Operations turned off the Linac beam in preparation for an access at 8:05 AM.

Operations reported at 8:08 AM that the Recycler had stashed four times during the midnight shift.

Operations loaded an energy conservation time line into the Time Line Generator at 8:18 AM.

Interlock experts reported at 8:27 AM that they would be conducting critical device controller (CDC) tests for the P1 and A1 transfer lines, and for Switchyard.

A Preaccelerator expert accessed the I- Source dome at 9:38 AM.

Operations established beam to NuMI and MiniBooNE at 1:40 PM.

Operations established beam to MTest at 2:15 PM.

Operations established beam to Pbar at 3:05 PM.

Experts began tuning the Accumulator stochastic cooling system at 3:28 PM

Pbar returned to stacking at 5:00 PM.

Operations disabled the MTest beam at 6:05 PM.

Friday Feb 20

The midnight shift began with Operations monitoring store 6816 (74.08E30) and the stash (109.7E10), with Pbar stacking (31.5E10), and with NuMI and MiniBooNE taking beam.

Operations received an E3 sump alarm at 2:10 AM.

Operations terminated store 6816 at 7:03 AM.

Experts began Tev studies at 7:31 AM.

The Plans for Friday and the Weekend

Access F sector and Pbar rings, put in a store, then return to stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/20/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory